Bubba:

Part A: Find a rational number that is between 5.2 and 5.5. Explain why it is rational. (2 points) Part B: Find an irrational number that is between 5.2 and 5.5. Explain why it is irrational. Include the decimal approximation of the irrational number to the nearest hundredth.

Vocaloid:

for the first part, a finite decimal like 5.3 should work since it can be expressed as the quotient of two real numbers

Vocaloid:

for part B, try to find a number that gives you a number between 5.2 and 5.3 when you take its square root

Bubba:

5.5?

Vocaloid:

nope

Bubba:

5.4

Vocaloid:

the square root of 5.5 is NOT between 5.2 and 5.3 keep trying.

Bubba:

Obviously math is my worst subject! lol

Vocaloid:

here's a hint: 5.2^2 = ? 5.5^2 = ?

Bubba:

do you mean "Not between 5.2 and 5.5?

Bubba:

5.3? This is so confusing

Vocaloid:

again, 5.2^2 = ? 5.5^2 = ?

Bubba:

5.2^2 =5.4? 5.5^2 = 5.8?

Vocaloid:

^ means exponent

Bubba:

ughhhhhhhhhhhhhhhhhh!

Vocaloid:

5.2^2 = 5.2 * 5.2 = ? 5.5^2 = 5.5 * 5.5 = ?

Bubba:

27.04 30.25

Vocaloid:

good so we can pick a number between those two (let's pick 29) so the square root of 29 is a possible answer for part B (this equals 5.39 rounded as your answer)